# Smallest Suffixient Sets as a Repetitiveness Measure

### v3

@@ 1 Introduction
In this paper we better characterize $\chi$ as a repetitiveness measure. First, we study how it behaves when the string undergoes updates, showing in particular that it grows by $O(1)$ when appending or prepending symbols, but that it can grow additively by $\Omega(\log n)$ upon arbitrary edit operations or rotations, and by $\Omega(\sqrt{n})$ when reversing the string. Second, we show that $\chi=O(r)$ on every string family, where $r$ is the number of equal-letter runs of the BWT of the string. We also show that there are string families where $\chi=o(v)$, where $v$ is the size of the smallest lexicographic parse [28] (an alternative to the size of the Lempel-Ziv parse [20], which behaves similarly). In particular, this holds on the Fibonacci strings, where we fully characterize the only 2 smallest suffixient sets of size 4, and further prove that $\chi\leq\sigma+2$ on all substrings of episturmian words over an alphabet of size $\sigma$. Since $v=O(r)$ on all string families, this settles $\chi$ as a strictly smaller measure than $r$, which is a more natural characterization than in terms of the reverse string. We also show that $\chi$ is incomparable with most “copy-paste” based measures [25], as there are families where it is strictly smaller and others where it is strictly larger than any of those measures.
@@ 4 Sensitivity of $\chi$ to String Operations
In this section we study how $\mathtt{sre}$ and $\chi$ behave in this respect.
@@ 7 Conclusions and Open Questions
Regarding edit operations, experiments suggest that $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)$ is $O(1)$ for all the string operations we considered. Proving a multiplicative constant for insertions would imply the existence of a constant for rotation and vice versa. It is also open whether $r=O(\chi\log\chi)$. If this were true —and provided that $\chi$ has $O(1)$ multiplicative sensitivity to string operations— it would imply that $r$ has $O(\log n)$ multiplicative sensitivity to these operations, making the already known lower bounds on multiplicative sensitivity [1, 14, 15] tight. If the conjecture were false, then $\chi$ could be considerably smaller than $r$ in some string families.
@@ Acknowledgements
We thank Davide Cenzato, Nicola Prezza, and Francisco Olivares for their code to compute smallest suffixient sets https://github.com/regindex/suffixient [5], which was helpful to propose and discard hypotheses on the behavior of $\chi$, for useful discussions on suffixient sets, and to verify our implementation of our Ukkonen’s extension that computes them online.

### v4

# Smallest Suffixient Sets: Effectiveness, Resilience, and Calculation

@@ 1 Introduction
In this paper we better characterize $\chi$ as a repetitiveness measure. First, we study how it behaves when the string undergoes updates, showing in particular that it grows by $O(1)$ when appending or prepending symbols, but that it can grow additively by $\Omega(\log n)$ upon arbitrary edit operations or rotations, and by $\Omega(\sqrt{n})$ when reversing the string. For this last operation, we show that $\chi$ cannot grow by a factor larger than $2$, however. Second, we show that $\chi=O(r)$ on every string family, where $r$ is the number of equal-letter runs of the BWT of the string. We also show that there are string families where $\chi=o(v)$, where $v$ is the size of the smallest lexicographic parse [28] (an alternative to the size of the Lempel-Ziv parse [20], which behaves similarly). In particular, this holds on the Fibonacci strings, where we fully characterize the only 2 smallest suffixient sets of size 4, and further prove that $\chi\leq\sigma+2$ on all substrings of episturmian words over an alphabet of size $\sigma$. Since $v=O(r)$ on all string families, this settles $\chi$ as a strictly smaller measure than $r$, which is a more natural characterization than in terms of the reverse string. We also show that $\chi$ is incomparable with most “copy-paste” based measures [25], as there are families where it is strictly smaller and others where it is strictly larger than any of those measures.
@@ 4 Sensitivity of $\chi$ to String Operations
In this section we study how $\mathtt{sre}$ and $\chi$ behave in this respect. We leave for the next section our results on the multiplicative sensitivity of $\chi$ to reversing the string.
@@ 4.4 Lower and upper bounds on the sensitivity of $\chi$
5 On the Multiplicative Sensitivity of $\mathtt{sre}$ to Reverse
We prove in this section that $\chi$ at most doubles when we reverse the string.
In the following, we denote the set of symbols preceding a substring $x$ (or left-extensions of $x$) of $w$ by $L_{x}(w)$, and we use $l_{x}(w)$ to denote an arbitrary element of $L_{x}(w)$ (we drop $w$ if it is clear from the context). For simplicity of analysis, we assume that $w$ contains a special symbol \$ at the beginning and at the end. This does not matter in asymptotic terms, as $\mathtt{sre}(w)=\mathtt{sre}(\text{\tt\textdollar}w\text{\tt\textdollar})\pm O(1)$, and similarly, $\mathtt{sre}(\text{\tt\textdollar}w^{R}\text{\tt\textdollar})=\mathtt{sre}(w^{R})\pm O(1)$. We do this to (i) ensure all right-maximal substrings of $w$ have at least two left-extensions, and hence, right-extensions have at least one left-extension, and (ii) to provide the necessary symmetry to prove Lemma 10 below.
Please note that left-extensions are not analogous to right-extensions. A right-extension is a string $xa$ such that $xb$ also occurs for some $b\neq a$, while a left-extension of $x$ is just any symbol preceding $x$. A left-extension can be unique, whereas right-extensions come (at least) in pairs. We will rely on the following observation.
Observation 5.1
If $xa$ is a super-maximal right-extension, then $L_{xa}\cap L_{xb}=\emptyset$ for any $b\neq a$.
Proof
Otherwise, $xa$ is not super-maximal, because there exists $c\in L_{xa}\cap L_{xb}$, for which $cxa\in\mathcal{F}_{w}$ and $cxb\in\mathcal{F}_{w}$, and thus $xa$ is a proper suffix of the right-extension cxacxa of the right-maximal string $cx$. ∎
Let us introduce a special tree of super-maximal extensions, which will be helpful to prove our result.
Definition 7
The prefix super-maximal right-extension tree (PSR tree) of $w$ is a compacted trie of all the super-maximal right-extensions of $w$. Since some of those strings can be prefixes of others, internal nodes of the trie may also represent strings, and their corresponding nodes are not compacted. Distinct leaves do not represent prefixes of each other.
The following result is key for our proof.
Lemma 9
Let $w$ be a string starting with a unique symbol \$. If $x$ is a right-maximal substring of $w$, and $xa$ is a super-maximal right-extension of $x$ in $w$, then $x^{R}$ is a right-maximal substring of $w^{R}$.
Proof
Because $xa$ is a right-extension in $w$, there exists $xb\in\mathcal{F}_{w}$ with $b\neq a$. Any symbol preceding $xa$ cannot precede $xb$, otherwise, $xa$ would not be super-maximal. Since $w$ starts with the unique symbol \$, if $x\neq\varepsilon$, then $xa$ and $xb$ have occurrences that are not prefixes of $w$, and hence, they are preceded by different symbols. It follows $x^{R}$ is right-maximal in $w^{R}$.∎
Lemma 10
It always holds that $\mathtt{sre}(w^{R})/2\leq\mathtt{sre}(w)\leq 2\cdot\mathtt{sre}(w^{R})$.
Proof
We define an injection $\lambda$ from the set of leaves and unary internal nodes (those with exactly one child) of the PSR tree $T(w)$ of $w$ into $\mathcal{S}_{r}(w^{R})$. We will identify tree nodes with the string they represent.
  • Leaf nodes: If $xa$ is (the supermaximal right-extension represented by) a leaf node, we let $\lambda(xa)=ux^{R}\cdot l_{xa}$ where $l_{xa}\in L_{xa}(w)$ (which exists because $w$ starts with \$, as discussed), and $u$ is chosen such that $ux^{R}\cdot l_{xa}\in\mathcal{S}_{r}(w^{R})$. At least one such $u$ must exist because $x^{R}\cdot l_{xa}\in\mathcal{F}_{w^{R}}$ and $x^{R}$ is right-maximal in $w^{R}$ by Lemma 9.
  • Unary internal nodes: If $xa$ a unary internal node with child $yb=xazb$, consider a right-extension xazcxazc of xazxaz, where $c\neq b$ (one must exist because xazxaz is right-maximal). We then let $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}\in\mathcal{S}_{r}(w^{R})$ where $l_{yc}\in L_{yc}(w)\subseteq L_{xa}(w)$. Note $L_{yc}$ is not empty because, as in the previous point, xazcxazc is a supermaximal right-extension, and $u$ exists because $z^{R}ax^{R}$ is right-maximal in $w^{R}$ by Lemma 9.
We now prove that $\lambda$ is an injection. Consider two leaves $xa$ and $yb$ of $T(w)$:
  1. 1.
    If $x[i]\neq y[i]$ for some $i$, then $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{yb}$, aligned from the right, differ at their position $\lambda(xa)[|\lambda(xa)|-i]\neq\lambda(yb)[|\lambda(yb)|-i]$, and therefore are different.
  2. 2.
    If $x$ is a proper prefix of $y$, then and $y$ has to start with $xc$ for some $c\neq a$ (otherwise $xa$ would be internal). Because $xa$ is a supermaximal right-extension, the set of left-extensions $L_{xa}$ and $L_{yb}\subseteq L_{xc}$ are disjoint by Observation 5.1. Therefore, the selected super-maximal extensions are of the form $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{xc}$ with $l_{xa}\neq l_{xc}$, thus, distinct.
  3. 3.
    Similarly, if $x=y$, the two super-maximal extensions $xa$ and $xb$ have disjoint sets of left-extensions $L_{xa}$ and $L_{xb}$, and hence $\lambda(xa)\neq\lambda(xb)$ because they end in different characters.
For a graphic depiction of these three cases see Figure 2.
In the other cases, one of the two nodes is a (unary) internal node. Assume, without loss of generality, that $|x|\leq|y|$. By the same argument of point 1 above, $\lambda(xa)$ cannot be equal to some $\lambda(yb)$ if $x$ is not a prefix of $y$. It follows that $xa$ is a unary internal PSR node that is an ancestor of $yb=xazb$:
  1. 4.
    Let $yb$ be the child of $xa$. Let $c\neq b$ be such that xazcxazc is a right-extension in $w$, and $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}$ with $l_{yc}\in L_{yc}$. Note that $\lambda(xa)$ differs from $\lambda(yb)$ because $\lambda(yb)$ finishes with some $l_{yb}\in L_{yb}$ and, by Observation 5.1, $L_{yb}\cap L_{yc}=\emptyset$.
  2. 5.
    If $yb$ is a farther descendant of $xa$, then xazxaz has children by $c$ and by some $d\neq c$ where xazdxazd prefixes $y$. The string xazdxazd is a supermaximal right-extension because there is a node for it in the tree, hence, by Observation 5.1, it holds $L_{xazc}\cap L_{xazd}=\emptyset$. As $L_{yb}\subseteq L_{xazd}$, it follows $L_{xazc}\cap L_{yb}=\emptyset$. Therefore, $\lambda(xa)$ and $\lambda(yb)$ end in different symbols.
For a graphic depiction of these two cases see Figure 3.
We are then injecting $k$ nodes into $\mathcal{S}_{r}(w^{R})$, where $k\geq\mathtt{sre}(w)/2$ (with equality when $T$ is a full binary tree). Thus, $\mathtt{sre}(w)/2\leq\mathtt{sre}(w^{R})$. Since the reversal operation is an involution, we have similarly that $\mathtt{sre}(w^{R})/2\leq\mathtt{sre}(w)$. ∎
[…]
Figure 2: Cases 1), 2), and 3) of Lemma 10, where $xa$ and $yb$ are leaves of the PSR tree. Note that the suffix $x^{R}\cdot l_{xa}$ and $x^{R}\cdot l_{yb}$ of $\lambda(xa)$ and $\lambda(yb)$ are not a suffix one of the other because the grayed symbols are different.
[…]
Figure 3: Case 4) and 5) of Lemma 10, where $xa$ is an internal node of the PSR tree, ybyb is the only child of $xa$, and $y^{\prime}d$ is a farther descendant of $xa$ (if it exists). Note xazcxazc is a right-extension, but not super-maximal (otherwise, $xa$ would not be unary). In this case, $\lambda(xa)$ is suffixed by $y^{R}\cdot l_{yc}$, whereas $\lambda(yb)$ and $\lambda(y^{\prime}d)$ are suffixed by $y^{R}\cdot l_{yb}$ and $y^{R}\cdot l_{y^{\prime}d}$, respectively. Note how $yb$ and all its descendants yield $\lambda(\cdot)$ values differing from $\lambda(xa)$ in their last symbol.
@@ 8 Conclusions and Open Questions
Regarding edit operations, experiments suggest that $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)$ is $O(1)$ for all the string operations we considered. We proved such a multiplicative constant for reversals. Proving the same for insertions would imply the existence of a constant for rotation and vice versa. It is also open whether $r=O(\chi\log\chi)$. If this were true —and provided that $\chi$ has $O(1)$ multiplicative sensitivity to string operations— it would imply that $r$ has $O(\log n)$ multiplicative sensitivity to these operations, making the already known lower bounds on multiplicative sensitivity [1, 14, 15] tight. If the conjecture were false, then $\chi$ could be considerably smaller than $r$ in some string families. In the same line, observe that Lemma 10 implies that, either $r(w)/r(w^{R})=\omega(\log n)$ and $r$ can be $\omega(\chi\log n)$, or $r(w)/r(w^{R})=O(\log n)$ and this known bound is tight.
@@ Acknowledgements
We thank Davide Cenzato, Nicola Prezza, and Francisco Olivares for their code to compute smallest suffixient sets https://github.com/regindex/suffixient [5], which was helpful to propose and discard hypotheses on the behavior of $\chi$, for useful discussions on suffixient sets, and to verify the implementation of our Ukkonen’s extension that computes them online.